### v2

# Collaborative Chain-of-Agents for Parametric-Retrieved Knowledge Synergy

Abstract
Retrieval-Augmented Generation (RAG) has emerged as a promising framework for enhancing the capabilities of Large Language Models (LLMs), especially in knowledge-intensive tasks. Despite its advantages, current RAG methods often struggle to fully exploit knowledge during generation. In particular, the synergy between the model’s internal parametric knowledge and external retrieved knowledge remains limited. Retrieved contents may sometimes mislead generation, while certain generated content can guide the model toward more accurate outputs. In this work, we propose Collaborative Chain-of-Agents, a framework designed to enhance explicitly synergy over both parametric and retrieved knowledge. Specifically, we first introduce CoCoA-zero, a multi-agent RAG framework that first performs conditional knowledge induction and then reasons answers. Building on this, we develop CoCoA, a long-chain training strategy that synthesizes extended multi-agent reasoning trajectories from CoCoA-zero to fine-tune the LLM. This strategy enhances the model’s capability to explicitly integrate and jointly leverage parametric and retrieved knowledge. Experiments results show that CoCoA-zero and CoCoA achieve superior performance on open-domain and multi-hop QA tasks.  [^1] Code available at https://github.com/liunian-Jay/CoCoA .
1 Introduction
Large Language Models (LLMs) (Achiam et al. 2023; Touvron et al. 2023) have demonstrated strong performance across a wide range of natural language tasks. However, the knowledge they rely on is embedded in their parameters and cannot be easily updated as new information emerges (Ji et al. 2023; He, Zhang, and Roth 2022). To address this limitation, the Retrieval Augmented Generation (RAG) framework introduces an external retrieval component that brings in external knowledge and integrates it into the input context of the LLMs. This design has led to notable improvements in various natural language processing applications (Gao et al. 2023; Lewis et al. 2020). Existing research has primarily aimed to improve two aspects of RAG: retrieving more relevant information during retrieval and better utilizing that information to guide generation during generation. Despite these efforts, most retrieval-augmented language models (RALMs) still emphasize external retrieval, while paying insufficient attention to the rich internal knowledge already encoded in model parameters. This internal knowledge is especially valuable for open-domain question answering, where many queries are factual and often already covered during pretraining.
Figure 1: Evaluation on 2WikiMultiHopQA, HotpotQA, and WebQuestions. The Merge method is a simple strategy we use to verify the collaboration of internal and external knowledge. It directly generates a passage and merges it into the retrieved passages as the context of the LLM.
Specifically, as the knowledge in LLM’s parameter becomes richer and the abilitiy of the LLM becomes stronger, sometimes answers with search information are not as good as direct answers. To validate the necessity of collaboratively synergizing internal (or parametric) and external (or retrieved) knowledge, we conducted experiments to compare performance. As shown in Fig. 1, across the three evaluation tasks, direct generation and GenRead (Yu et al. 2022) (explicitly generated content) sometimes shows stronger performance. Also, we conduct a test experiment,“Merge”, that explicitly integrates internal and external knowledge by combining retrieved passages with internally generated passages as the final context, as shown in Fig. 1. “Merge” often achieves better results than both direct generation and RAG approaches, demonstrating the potential of internal and external knowledge collaboration. However, its improvements are not consistent across all datasets, indicating the need for more sophisticated integration methods.
Figure 2: Illustration of the CoCoA framework. The top part is CoCoA-zero, a multi-agent collaboration framework. It integrates internal and external knowledge in a collaborative manner by first performing knowledge induction and then making decisions. The bottom part is the training strategy, which is based on CoCoA-zero and combines the trajectories of different agents into long chains to train and enhance the integration ability of the LLM.
Existing methods solve the problem of knowledge collaboration through RAG pipeline optimization. Some approaches alleviate this through workflow or multi-module collaboration. For example, SURE (Kim et al. 2024a) generates multiple candidate answers and verifies them one by one to ensure reliability. CON (Yu et al. 2023) mitigates external noise by adding a processing chain. AstuteRAG (Wang et al. 2024a) integrates reliable information iteratively. There are also some approaches that solve the problem of knowledge collaboration through enhanced training of the LLM. For instance, RAFT (Zhang et al. 2024) employs anti-noise training to enable the model to effectively utilize internal knowledge when external documents contain noise, while Self-RAG (Asai et al. 2023) learns to determine whether retrieval is needed in advance, thereby avoiding harmful content before retrieval. Despite these efforts, existing work still has notable limitations. On the one hand, some pipeline optimization methods tend to lose effectiveness as LLMs become more capable. On the other hand, some fine-tuning methods cannot fully exploit internal knowledge effectively.
To address the above challenges, we introduce CoCoA, which consists of a multi-agent reasoning framework and a training strategy that combines multi-agent trajectories into long chains to enhance LLM performance. Specifically, we first introduce CoCoA-zero, which features three complementary agents: one for extracting pre-trained knowledge, one for retrieving external data, and one for reasoning over both to make optimal decisions. This not only enables explicit construction of decoupled internal and external knowledge, but also provides collaborative reasoning traces for the training, particularly the agent’s ability to synthesize information and make context-aware decisions. Based on CoCoA-zero, we further introduce an end-to-end training strategy for CoCoA, which significantly improves performance on knowledge-intensive tasks by fusing the collaborative capabilities of multi- agents into one model.
In general, our contributions are summarized as follows:
  • We investigate the challenge of parametric–retrieved knowledge collaboration and introduce CoCoA-zero, a multi-agent reasoning framework that coordinates parametric and retrieved knowledge for improved generation.
  • We develop a training paradigm for CoCoA, which distills multi-agent reasoning into long-chain, enabling LLMs to better exploit internal and external knowledge.
  • Extensive experiments demonstrate CoCoA’s effectiveness, offering insights for inference-time scaling and multi-agent training on knowledge-intensive tasks.
2 Methodology
In this section, we present CoCoA-zero and CoCoA, as illustrated in Fig. 2. We first describe the multi-agent framework, CoCoA-zero, followed by the long-chain training strategy for CoCoA. The algorithm is shown in Algorithm 1.
2.1 Preliminaries
We formalize the standard Retrieval-Augmented Generation framework. Given a query $q$ and a corpus $\mathcal{D}$, the RAG system retrieves $k$ relevant passages $C=\{c_{1},c_{2},\cdots,c_{k}\}\subset\mathcal{D}$ and generates an answer $\hat{a}$ based on the combined input. This process follows a retrieve-then-generate paradigm and can be formulated as:
$\begin{split}C&=\mathcal{R}(q,\mathcal{D},k),\\ \hat{a}&=\mathcal{G}(\mathcal{P}(q,C)),\end{split}$ (1)
where $\mathcal{R}$ is the retriever, $\mathcal{P}$ is the prompt constructor that formats $q$ and $C$, and $\mathcal{G}$ is the generator (e.g., a LLM) that predicts the final answer $\hat{a}$.
2.2 Two-stage RAG Framework: CoCoA-zero
In this section, we present our multi-agent RAG framework, CoCoA-zero, which also functions as the data synthesis pipeline for CoCoA. As shown in Fig. 2, Stage I (§ 2.2) employs two specialized agents to induce knowledge from parameters and retrieval, while Stage II (§ 2.2) introduces a agent to synthesize them for high-level decision-making.
Stage I: Knowledge Induction.
It is challenging to extract implicit knowledge solely from the model’s internal knowledge or retrieved passages. Inspired by GenRead (Yu et al. 2022) and SURE (Kim et al. 2024a), we design two dedicated agents for knowledge induction. Each agent first generates an answer to the question and then summarizes knowledge based on that answer.
Induction of Internal Knowledge. Directly allowing the model to explicitly generate its own internal knowledge is difficult to control and will inevitably result in sparse or inconsistent knowledge being generated. Following SURE (Kim et al. 2024a), we introduce conditional induction. Specifically, the Internal Knowledge Agent samples a candidate $a_{\text{in}}$ from the LLM based on the question:
$a_{\text{in}}=G(\mathcal{P}(q))$ (2)
Next, we prompt the LLM to generate a knowledge passage $s_{\text{in}}$ conditioned on $q$ and $a_{\text{in}}$, which reflects the model’s internal understanding:
$s_{\text{in}}=G(\mathcal{P}(q,a_{\text{in}})).$ (3)
Induction of External Knowledge. For retrieved passages, the External Knowledge Agent follows a similar procedure. Specially, it first retrieve some passages $C=\{c_{1},c_{2},\cdots,c_{k}\}$ from the corpus $\mathcal{D}$. Conditioned on both $q$ and $C$, it produces a second candidate $a_{\text{ext}}$:
$a_{\text{ex}}=G(\mathcal{P}(q,C))$ (4)
Then, conditioned on $q$, $a_{\text{ex}}$ and $C$, the agent induces the external knowledge passage $s_{\text{ex}}$, :
$s_{\text{ex}}=G(\mathcal{P}(q,a_{\text{ex}},C)).$ (5)
The conditional knowledge induction framework makes implicit knowledge explicit and controllable, improving the model’s ability to express knowledge and providing a solid basis for high-level decision-making in the next stage.
Stage II: High-level Decision Making.
Building on the candidate answers and inductive knowledge obtained in Stage I, the second stage leverages the LLM’s reasoning ability to perform high-level decision making.
Specifically, the Decision-Making Agent adopts COT (Wei et al. 2022) reasoning over the internal and external candidate answers and their corresponding knowledge. It will be prompted with all five components (questions, internal and external candidate answers and their corresponding inductive knowledge) and generate the final answer $\hat{a}$ through COT.
$cot_{\text{a}},\hat{a}=G(\mathcal{P}_{\text{cot}}(q,s_{\text{in}},a_{\text{in}},s_{\text{ex}},a_{\text{ex}}))$ (6)
Here, $cot_{\text{a}}$ denotes the reasoning path that drives explicitly decision-making and guides final answer generation.
The model thereby functions as a high-level aggregator, reinforcing potentially consistent beliefs and resolving potential conflicts between internal beliefs and retrieved evidence. By explicitly modeling and comparing knowledge before committing to an answer, our framework improves the transparency and robustness of the decision process.
2.3 Collaborative Chain-of-Agents Training
Although multi-agent collaboration for internal and external knowledge coordination is simple and effective, how to achieve global optimization across multi-agents remains non-trivial.
To this end, we propose the Collaborative Chain-of-Agents training strategy, which aims to optimize multi-agent collaboration end to end by supervising the LLM on long-form reasoning trajectories. These trajectories are synthesized from the multi-agent pipeline CoCoA-zero (§ 2.2) and reflect the full reasoning process that integrates both parametric and retrieved knowledge.
Figure 3: Illustration of the training for CoCoA.
Supervised Fine-Tuning.
The CoCoA-zero framework is designed to: (1) control the direction of knowledge generation via conditional induction, (2) decouple internal and external knowledge through parallel reasoning, and (3) integrate both sources via Chain-of-Thought decision making.
To supervise the model to achieve explicit and collaborative knowledge integration, we synthesize training samples by concatenating the intermediate results produced by CoCoA-zero into a single long-form response. Specifically, given a question $q$ and a set of retrieved documents $C$, we integrate the intermediate results from the CoCoA-zero (i.e., internal induction $s_{\text{in}}$, external induction $s_{\text{ex}}$, the CoT reasoning trace $cot_{\text{a}}$ during integration and the final answer $\hat{a}$ ) into a long response $y$ and promote the evolution of model capabilities through the following supervision objectives:
$\mathcal{L}_{\text{SFT}}=-\mathbb{E}_{(x,y)\sim\mathcal{D}}\big{[}\log P_{\theta}(s_{\text{in}},s_{\text{ex}},cot_{\text{a}},\hat{a}\mid q,d)\big{]}.$ (7)
This training explicitly exposes the model to long collaborative samples, where the target outputs are synthesized based on the outputs of CoCoA-zero. Through end-to-end training, multiple agents can influence and enhance each other’s capabilities. Moreover, the noise introduced by intermediate agents becomes negligible, as it contributes to the overall robustness of the training process.
Algorithm 1 CoCoA: Example of one sample
Input: Query $q$, corpus $\mathcal{D}$, hyperparameters $k$
Output: Final answer $\hat{a}$ or training sample $y$
1:CoCoA-zero:
[…]
2:if Supervised Fine-tuning then
[…]
4:  Update model with $\mathcal{L}_{\text{SFT}}$ in Eq. 7.
5:end if
6:if DPO Training then
7:  $y^{-}\leftarrow G(\mathcal{P}_{\text{ZS}}(q,C))$
[…]
9:  Update model with $\mathcal{L}_{\text{DPO}}$ in Eq. 8
10:end if
11:return $\hat{a}$ or the trained model CoCoA
Preference optimization.
[…]
$\mathcal{L}_{\text{DPO}}(\pi_{\theta})=$ $-\mathbb{E}_{(x,y^{+},y^{-})\sim\mathcal{D}}\big{[}\log\sigma\big{(}\beta\cdot\log\pi_{\theta}(y^{+}|x)$ (8)
$-\beta\cdot\log\pi_{\theta}(y^{-}|x)\big{)}+\alpha\cdot\big{(}-\log\pi_{\theta}(y^{+}|x)\big{)}\big{]}$
where $\pi_{\theta}(y|x)$ denotes the unnormalized log-probability of response $y$ under the model $\theta$.
The CoCoA traing thus bridges symbolic multi-agent collaboration and end-to-end generation, enabling the model to internalize structured reasoning through supervision.
2WikiMQA HotpotQA WebQuestions TriviaQA
Method EM F1 Avg EM F1 Avg EM F1 Avg EM F1 Avg
Llama-3.1-Instruct Train-free & w/o retrieval
Llama-3.1-8B 27.60 28.35 27.98 24.00 27.09 25.54 40.11 39.98 40.04 62.87 64.17 63.52
8B+COT 23.80 26.55 25.28 26.20 32.26 29.23 38.04 39.43 38.73 64.90 66.98 65.94
8B+GenRead 24.00 23.92 23.96 29.20 31.15 30.18 29.53 29.67 29.60 54.12 54.29 54.21
Llama-3.1-70B 33.80 33.43 33.62 37.00 37.89 37.45 44.83 43.92 44.38 77.89 78.93 78.81
Llama-3.1-Instruct Train-free & w/ retrieval
8B+StandardRAG 26.80 25.07 25.94 31.40 34.16 32.78 37.65 37.32 37.49 66.83 67.16 66.99
8B+COT 22.40 25.25 23.83 32.40 38.71 35.55 35.73 36.17 35.95 65.85 67.54 66.69
8B+CON 19.00 21.32 20.16 32.80 38.67 35.73 34.40 38.05 36.22 65.64 66.82 66.23
8B+SURE 18.40 21.32 19.86 32.00 37.26 34.63 32.48 39.01 35.75 63.14 62.91 63.02
CoCoA-zero-8B 31.40 31.92 31.66 37.40 41.20 39.30 43.11 39.13 41.12 70.73 69.99 70.36
Llama-3.1-70B 22.00 23.12 22.56 35.20 38.03 36.61 39.76 39.05 39.41 70.97 71.44 71.20
RALM w/ retrieval & w/ Training
Self-RAG 7B 37.40 17.93 27.66 33.40 20.57 26.99 44.64 25.75 35.19 66.30 37.27 51.78
Self-RAG 13B 38.80 22.61 30.71 35.40 21.64 28.52 45.87 25.31 35.59 68.74 38.22 53.48
DeepSeek-R1-8B 36.80 25.79 31.30 35.00 32.66 33.83 44.34 31.87 38.11 65.62 58.07 61.84
InstructRAG-8B 36.40 39.40 37.90 - - - - - - 70.90 65.40 68.15
CoCoA-SFT-8B 41.00 36.87 38.94 39.40 46.31 42.86 42.96 41.32 42.14 70.72 70.39 70.55
CoCoA-DPO-8B 42.00 40.58 41.29 39.00 43.39 41.20 44.83 42.21 43.52 71.52 70.42 70.97
Table 1: EM/F1 of different methods experimented on four datasets. The best and second best scores are highlighted in bold and underlined, respectively. Italics mark a boundary, not for comparison. represents the Out-of-Distribution evaluation dataset.
2.4 Optimization Analysis
We compare independent training and CoCoA training under a simplified two-step setting involving pre-generation processing followed by answer generation:
$\mathcal{L}{\text{indep}}=-\log P_{\theta}(s|x,d)-\log P_{\phi}(\hat{a}|s)$ (9)
$\mathcal{L}{\text{chain}}=-\log P_{\theta}(s|x,d)-\log P_{\theta}(\hat{a}|x,d,s)$ (10)
Gradient comparison:
$\frac{\partial\mathcal{L}{\text{chain}}}{\partial\theta}=\frac{\partial\mathcal{L}{\text{indep}}}{\partial\theta}+\Delta_{g}$ (11)
where $\Delta_{g}:=\frac{\partial}{\partial\theta}\bigl{[}-\log P_{\theta}(\hat{a}|x,s,d)\bigr{]}$. $\Delta_{g}$ captures feedback from the answer to the pre-processing, which is absent in independent training. Chain training is a special type of multi-task learning that helps to break out of local optimization. The experimental results are in Section 3.6, and detailed derivations are in Appendix D.
3 Experiments
In this section, we report our experiments results, and provide a analysis of them.
3.1 Implementation Details
Training Data
We sample subsets from the training sets of HotpotQA (Ho et al. 2020a), 2WikiMultiHopQA (Ho et al. 2020b) and WebQuestions (Berant et al. 2013), then synthesize data using the CoCoA-zero and filter them based on gold answers. This results in 6.8k filtered samples for SFT. For DPO, we select 1151 samples, which are the ones that are answered incorrectly by zero-shot but correctly by the CoCoA-zero framework. For each sample, we gather 5 relevant passages using CONTRIEVER (Izacard et al. 2021).
Training Details
We fine-tune LLaMA3.1-8B with LoRA (r=16, $\alpha$=16, dropout=0.05). During SFT, we train for 5 epochs with a learning rate of 3e-5. For DPO, we used $\beta$=0.2 and $\alpha$=0.2 (RPO), with a learning rate of 5e-6. All experiments are conducted on a single A100 GPU.
Inference Details
During inference, we use Contriever (Izacard et al. 2021) as the retriever and set k to 5. For all datasets, we use 21M English Wikipedia (Karpukhin et al. 2020) dump as the source passages for the retrieval. Prompts for the experiments can be found in Appendix F.
3.2 Datasets and Evaluation Metrics
Eval Datasets To evaluate the effectiveness and generalization of CoCoA, we conduct experiments on two open-domain question answering datasets: WebQuestions (Berant et al. 2013), and TriviaQA (Joshi et al. 2017), as well as two multi-hop question answering benchmarks: HotpotQA (Ho et al. 2020a) and 2WikiMultiHopQA (Ho et al. 2020b). Dataset statistics are summarized in Table 2, and further details are provided in Appendix A.
Task Type Datasets # Samples
Multi-HopQA 2WikiMultiHopQA 500
HotpotQA 500
OpenQA WebQuestions 2032
TriviaQA 11313
Table 2: Description of tasks and evaluation datasets.
Evaluation Metrics
We report both exact match (EM) and F1 scores. Following Asai et al. (2023); Mallen et al. (2022), we adopt a non-strict EM metric that deems a prediction correct if it contains the gold answer. F1 measures token-level overlap between the predicted and gold answers. In our setting, longer responses often yield higher EM scores due to increased coverage, but may reduce F1 scores by introducing irrelevant content. Thus, considering both metrics provides a more balanced evaluation.
3.3 Baselines
We selected several of the most representative methods for comparison. 1) StandardRAG, which is the most classic “retrieve-then-read” paradigm. 2) Chain-Of-Thought (Wei et al. 2022): Uses CoT prompting to generate reasoning steps before producing the final answer. 3) Chain-Of-Note (Yu et al. 2023): Refines the retrieved passages prior to answering. 4) GenRead (Yu et al. 2022): Generates self-contained intermediate context to answer, effectively replacing retrieval with generation. 5) SURE (Kim et al. 2024a): Conditional summarization followed by multiple validation. 6) Self-RAG (Asai et al. 2023): Employs adaptive retrieval and self-reflection to decide when and how to use external knowledge. 7) DeepSeek-R1-Distill-8B (Guo et al. 2025): A distilled LLaMA-8B model released by DeepSeek-R1, trained on reasoning data. 8) InstructRAG (Wei, Chen, and Meng 2024): Denoising training using self-synthesized data. All retrieval-based methods use top-5 passages. Other experimental settings follow those reported in the original papers. Other experimental settings are shown in the Appendix B.
3.4 Main Results
Experimental results are presented in Table 1, and we summarize the key findings as follows:
Retrieval vs. non-retrieval methods
On WikiMQA and WebQuestions, direct generation performs better, while retrieval methods excel on other tasks. This demonstrates that retrieved knowledge and parametric knowledge each have their own strengths and weaknesses in different scenarios.
RAG without training.
The improvements of some process optimization methods are decreasing compared to standardRAG. We speculate that this is because current LLMs are becoming more powerful enough to make good use of external knowledge. CoCoA-zero improves the average EM and F1 of all tasks by 4.99% and 4.64% respectively, while other train-free methods show little effect. These results suggest that current QA tasks should place greater emphasis on leveraging the model’s rich internal knowledge.
Superiority and Generalization of CoCoA.
Our CoCoA methods achieve state-of-the-art performance across almost all datasets. In particular, CoCoA improves the EM and F1 of 2WikiMultiHopQA tasks by 15.2% and 15.51% respectively. Moreover, despite being trained with limited data, CoCoA also performed well on other out-of-distribution datasets like TriviaQA, demonstrating its robustness.
Reasoning Distillation vs. CoCoA Training.
DeepSeek-R1-8B, trained on distilled reasoning data, outperforms the undistilled StandardRAG. CoCoA, distilled with multi-agent self-synthesis on knowledge-intensive tasks, further surpasses DeepSeek-R1-8B. We speculate this is because logical reasoning and knowledge-intensive tasks differ, and CoCoA can better leverage knowledge. This suggests that explicitly leveraging key internal and external knowledge can be more effective than chain-of-thought reasoning.
Benefit of DPO Training.
Comparing our supervised and DPO variants, DPO training yields improvements across several datasets. This suggests that contrastive preference learning can help the model better align to the collaborative responses of multi-agents. However, it may also lead to performance degradation due to the quality of training data.
Figure 4: Performance varies with the number of documents: 2WikiMQA (left), HotpotQA (middle), WebQuestions (right).
3.5 Ablation Study I: Different Modules
To better understand the contribution of each module in CoCoA-zero, we conduct an ablation study by selectively removing internal/external induction and the reasoning.
As shown in Table 3, removing internal induction significantly degrades performance, especially by 8.4% on 2WikiMQA. This shows the importance of leveraging parameterized knowledge in scenarios such as 2WikiMultiHopQA where the LLM itself can answer well. Similarly, excluding external induction also leads to a noticeable performance drop across all datasets, highlighting the complementary role of retrieved knowledge. Moreover, disabling the reasoning mechanism in decision making results in a consistent decrease, suggesting that reasoning over both knowledge contributes to deeper understanding.
To further validate the effectiveness of multi-agent collaboration, we introduce a zero-shot variant using a single agent. Its performance is much lower than CoCoA-zero, which confirms the necessity of using multi-agent roles to coordinate between internal and external knowledge.
Overall, these results confirm the effectiveness of our multi-agent collaboration design, where each component plays a non-trivial role in achieving optimal performance.
Method 2WikiMQA HotpotQA WebQuestions
CoCoA-zero 31.66 39.30 41.12
w/o Internal 23.26 ($\downarrow$ 8.40) 36.56 ($\downarrow$ 2.74) 39.10 ($\downarrow$ 2.02)
w/o External 28.97 ($\downarrow$ 2.69) 30.96 ($\downarrow$ 8.34) 38.97 ($\downarrow$ 2.15)
w/o Think 30.38 ($\downarrow$ 1.28) 37.17 ($\downarrow$ 2.13) 39.75 ($\downarrow$ 1.37)
Zero-Shot 18.55 ($\downarrow$ 13.11) 35.01 ($\downarrow$ 4.29) 35.38 ($\downarrow$ 5.74)
Standard 25.94 ($\downarrow$ 5.72) 32.78 ($\downarrow$ 6.52) 37.49 ($\downarrow$ 3.63)
Table 3: Ablation study on knowledge induction and decision-making. The zero-shot variant (§ 2.3) is also included. We use the average of EM and F1 for fair evaluation.
3.6 Ablation Study II: Training Strategies
To evaluate the effectiveness of our training strategy for CoCoA, we conduct an ablation study comparing different training configurations on the LLaMA3.1-8B model. As shown in Table 4, $\textit{Long-DPO}_{8B}$ achieves the best overall performance, confirming the benefit of aligning long-form outputs via long-chain optimization.
The $\textit{Short-SFT}_{8B\times 3}$ variant, where each task segment is trained on a separate model, shows clear degradation in performance, especially on 2WikiMultiHopQA. This indicates that separating induction and reasoning capabilities into isolated modules weakens the model’s ability to holistically integrate information across steps. The $\textit{Short-SFT}_{8B}$ variant, which combines three instruction capabilities into a single model but retains short-form generation, performs better than $\textit{Short-SFT}_{8B\times 3}$ but still falls behind our approaches. This shows that simply merging instructions is slightly less performant than our long chain consolidation.
Our training strategy for CoCoA, represented by $\textit{Long-DPO}_{8B}$ and $\textit{Long-SFT}_{8B}$ variants, explicitly modeled multi-agent collaboration as a unified long-form output. The superior performance of these models underscores the advantage of training models to generate cohesive and contextually rich responses rather than fragmented predictions. This, to a certain extent, provides new perspectives for the expansion of knowledge-intensive long chains.
Method 2Wiki HotpotQA WebQ Average
$\text{Long-DPO}_{8B}$ 41.29 41.20 43.52 42.00
$\text{Long-SFT}_{8B}$ 38.94 42.86 42.14 41.31
$\text{Short-SFT}_{8B}$ 33.91 40.04 40.13 38.03
$\text{Short-SFT}_{8B\times 3}$ 28.31 40.58 39.84 36.24
Table 4: Ablation study of the training strategy for CoCoA. For fairness, the average of EM and F1 is used as the metric.
3.7 When the Number of K Changes
In order to better explore the robustness of our CoCoA with respect to the number of documents, we set $K$ to vary in the interval [1, 3, 5, 10, 15, 20]. The results are shown in Fig. 4. Overall, our method outperforms StandardRAG across different values of $K$. Moreover, our method achieves stronger performance than StandardRAG when given less context. We speculate that this is because our model can better utilize internal knowledge, especially when given less information. However, our advantage decreases when the number of documents is too large. We speculate that this is due to the long context bottleneck of the model.
In summary, our method demonstrates strong robustness across different context sizes and provides a practical solution in settings with limited external information or constrained retrieval capacity.
3.8 Performance of Different Model Sizes
To verify the performance difference of CoCoA-zero under different model sizes, we conducted experiments on performance changes of different model sizes. As shown in Fig. 5, the larger the LLM, the better the performance of CoCoA-zero, and it far exceeds standardRAG. This shows that larger models better support our collaboration and highlights the importance of internal knowledge in stronger LLMs: the more powerful the LLM, the more it should leverage its internal knowledge for question answering.
Figure 5: Illustration of accuracy changes at different model sizes, with Avg(EM,F1) as the metric.
3.9 Training Generalization to Non-QA Tasks
To further evaluate the generalization ability of CoCoA, we test its performance on fact verification and multiple-choice tasks. As shown in Figure  6, our training did not reduce the performance of these tasks compared to standard RAG. In fact, in some cases, we even observed a slight improvement. One explanation is that our training strategy encourages collaborative output that leverages the capabilities of the LLM, rather than injecting knowledge directly, and thus possesses a certain degree of universality.
Figure 6: Illustration of accuracy changes when transferring to non-QA tasks, with accuracy as the metric.
4 Related Works
4.1 Retrieval-augmented Generation
In recent years, to address outdated knowledge and hallucination of LLM, RAG has been introduced (Fan et al. 2024; Gao et al. 2023), and many efforts have been made in two aspects: “how to retrieve more relevant information” including retriever fine-tuning (Nian et al. 2024) and query optimization (Ma et al. 2023; Wang, Yang, and Wei 2023; Wang et al. 2024b) and “how to better use the retrieved information” including domain fine-tuning (Wang et al. 2024c; Zhang et al. 2024; Yue et al. 2025; Xia et al. 2025) and controlled decoding strategies (Shi et al. 2023). Our CoCoA falls into the second category: better utilization of knowledge.
4.2 RAG Pipeline Optimization
Pipeline optimization usually adds pre-generation processing, retrieval intent identification, or optimizes the pipeline as a whole. For example, Glass et al. (2022); Kim and Lee (2024) and Yu et al. (2023) introduce reranking and refinement steps before generation, mitigating the impact of noisy retrieved passages. SKR (Wang et al. 2023) and UAR (Cheng et al. 2024) avoid unnecessary retrieval by adding retrieval intent identification processes before generation. SURE(Kim et al. 2024a) first generates multiple candidate answers and performs conditional summary verification based on the candidate answers, allowing LLMs to focus on specific contexts. However, these methods either rely too much on retrieved content or fail to combine internal and external knowledge, which can limit performance.
4.3 RALM Enhancement
Retrieved-Augmented Language Model(RALM) enhancement is usually achieved by adjusting the LLM to achieve effective use of the information. One common approach is to train the LLM itself. For example, RAFT (Zhang et al. 2024) and InstructRAG (Wei, Chen, and Meng 2024) improve the model’s ability to resist noise in external context by introducing noise resistance training. REAR (Wang et al. 2024c) achieves the model’s trade-off between external context and internal knowledge by training the model’s relevance-guided generation capabilities. Self-RAG (Asai et al. 2023) trains LLMs to decide whether to perform retrieval and to improve their self-reflection ability. Another approach involves guiding the decoding (Shi et al. 2023; Kim et al. 2024b). For instance, CAD (Shi et al. 2023) enforces absolute trust in retrieved information by using contrastive decoding under the assumption that external information is fully correct. However, both approaches tend to underutilize the model’s internal knowledge, which may constrain the quality and informativeness of its responses.
5 Conclusion
We investigate the challenge of parametric–retrieved knowledge collaboration and introduce CoCoA, a retrieval-augmented generation framework that improves LLM performance. By leveraging a two-stage multi-agent pipeline, CoCoA-zero collaborates internal and external knowledge and provides self-synthesized supervisory signals. Based on a long-chain training strategy, CoCoA delivers strong results on QA tasks, demonstrating its effectiveness and offering insights into long-chain reasoning and collaborative agent training for knowledge-intensive applications.
Appendix A Dataset
Here, we introduce in detail the datasets we used, which are four datasets on four tasks.
2WikiMultiHopQA (Ho et al. 2020b) and HotpotQA (Ho et al. 2020a): Both datasets are multi-hop question answering datasets based on Wikipedia. Considering the limitation of experimental cost, we used the sub-sampling set published by Trivedi et al. (2022); Kim et al. (2024a), which is obtained by extracting 500 questions from the validation set of each dataset.
WebQuestions (Berant et al. 2013): Constructed from questions posed by the Google Suggest API, where the answers are specific entities listed in Freebase.
TriviaQA (Joshi et al. 2017): A compilation of trivia questions paired with answers, both originally pulled from online sources.
Training Data
We sampled subsets from the training sets of HotpotQA (Ho et al. 2020a), 2WikiMultiHopQA (Ho et al. 2020b) and WebQuestions (Berant et al. 2013), then used the CoCoA-zero framework to synthesize data and filtered them with gold answers. Finally, we selected 6.8k filtered samples, including 3k, 3k, and 0.8k from the three datasets, respectively. For the DPO training data, we screen out 1151 samples, which are the ones that are answered incorrectly by zero-shot but correctly by the CoCoA-zero. For each sample, we gathered 5 relevant passages using the most common retriever Contriever (Izacard et al. 2021).
Appendix B Baseline Setting
We followed the original settings for almost all experiments. For baselines requiring training, we directly used their weights. Note that InstructRAG directly generates long rationales, the first half of which consists mostly of analysis and citations of the document, resulting in a non-strictly high EM score and a low F1 score. For a fair comparison, we used Qwen2.5-3B to perform answer segmentation to evaluate.
Appendix C Training Details
We fine-tune LLaMA3.1-8B with LoRA (r=16, $\alpha$=16, dropout=0.05) on a maximum input length of 2048. LoRA is applied to attention projection layers. During SFT, we trained for 5 epochs with a batch size of 1, gradient accumulation of 4, and a learning rate of 3e-5. For DPO, a $\beta$ value of 0.2 is applied, using a sigmoid loss function, while RPO is configured with an $\alpha$ value of 0.2. The learning rate was set to 5e-6 and other settings are the same as SFT. During inference, we use the vllm (Kwon et al. 2023) accelerated inference framework, and to ensure repeatability, we set the temperature to 0.0. All experiments are conducted on a single A100 GPU with 80GB or 40GB memory.
Method 2WikiMultiHopQA HotpotQA WebQuestions
EM F1 Avg EM F1 Avg EM F1 Avg
CoCoA-zero 31.40 31.92 31.66 37.40 41.20 39.30 43.11 39.13 41.12
w/o Thinking 30.00 30.76 30.38 36.00 38.34 37.17 39.17 40.32 39.75
w/o Internal 22.60 23.93 23.26 34.00 39.11 36.56 40.01 38.20 39.10
w/o External 28.40 29.53 28.97 30.00 31.92 30.96 39.81 38.13 38.97
Zero-Shot 17.60 19.51 18.55 33.20 36.81 35.01 34.45 36.31 35.38
Standard RAG 26.80 25.07 25.94 31.40 34.16 32.78 37.65 37.32 37.49
Table 5: Ablation study of internal/external induction and reasoning in decision making. In addition, a zero-shot method for explicit internal and external knowledge integration is added for comparison. For simplicity and fairness, the average of EM and F1 is used as the metric.
Method 2WikiMultiHopQA HotpotQA WebQuestions
EM F1 Avg EM F1 Avg EM F1 Avg
$\text{Long-DPO}_{8B}$ 42.00 40.58 41.29 39.00 43.39 41.20 44.83 42.21 43.52
$\text{Long-SFT}_{8B}$ 41.00 36.87 38.94 39.40 46.31 42.86 42.96 41.32 42.14
$\text{Short-SFT}_{8B}$ 28.60 28.03 28.31 39.00 42.15 40.58 41.19 38.48 39.84
$\text{Short-SFT}_{8B\times 3}$ 35.00 32.81 33.91 37.60 42.48 40.04 41.29 38.96 40.13
Table 6: Ablation study of the training strategy for CoCoA. For simplicity and fairness, the average of EM and F1 is used as the metric
Appendix D Optimization Analysis
We analyze the difference between independent training and long chain training in terms of the form of loss. We simplify the steps in this analysis, i.e., there are only two steps in the chain, pre-generation processing first and then answering.
When the two agents optimize independently, the loss takes the following form:
$\mathcal{L}_{\text{indep}}$ $=-\log P_{\theta}\!\bigl{(}s\mid x,d\bigr{)}\;-\;\log P_{\phi}\!\bigl{(}\hat{a}\mid s\bigr{)}.$ (12)
Here, $\theta$ and $\theta^{\prime}$ are optimized independently.
When two agents use long chain optimization, the loss is as follows:
$\mathcal{L}_{\text{chain}}$ $=-\log P_{\theta}\!\bigl{(}s,\hat{a}\mid x,d\bigr{)}$ (13)
$=-\log P_{\theta}\!\bigl{(}s\mid x,d\bigr{)}\;-\;\log P_{\theta}\!\bigl{(}\hat{a}\mid x,d,s\bigr{)}.$
Gradient propagation:
The gradient of the first term in Eq. (12) is,
$\frac{\partial}{\partial\theta}\mathcal{L}_{\text{indep}}$ $=\frac{\partial}{\partial\theta}\ \bigl{[}-\log P_{\theta}(s\mid x,d)\bigr{]}$ (14)
The gradient of the Eq. (13) is,
$\frac{\partial}{\partial\theta}\mathcal{L}_{\text{chain}}$ $=\frac{\partial}{\partial\theta}\bigl{[}-\log P_{\theta}(s\mid x,d)\bigr{]}+\frac{\partial}{\partial\theta}\bigl{[}-\log P_{\theta}(\hat{a}\mid x,s,d)\bigr{]}$ (15)
[…]
Here, $\Delta_{g}:=\frac{\partial}{\partial\theta}\bigl{[}-\log P_{\theta}(\hat{a}|x,s,d)\bigr{]}$ is the additional gradient that the answer‑loss naturally back‑propagates to the pre-processing parameters when the same network $\theta$ produces both tokens. In the independent setting $\Delta_{g}=0$ by construction, so the preprocessor never “hears” whether the answer is correct, which is not conducive to the consistency of the response. The chain objective restores this missing credit assignment signal, thus performing a special kind of multi-task learning on both stages, optimizing them instead of each in isolation, potentially helping to escape from local optimal solutions.
Appendix E Full Results
We supplemented the detailed results of the ablation experiment as shown in Table 5 and Table 6.
Appendix F Prompt Templates
All the prompt templates used by our proposed CoCoA are shown in Table 9 and Table 8. And special instructions are added to section 3.9 corresponding to different tasks as shown in Table 7.
Task Task Instruction
ARC-C Given four answer candidates, A, B, C and D, choose the best answer choice. Please answer with the capitalized alphabet only, without adding any extra phrase or period. Do not exceed one word.
PubHealth Is the following statement correct or not? Say true if it’s correct; otherwise say false. Don’t capitalize or add periods, just say ”true” or ”false”. Do not exceed one word.
Table 7: Full list of instructions used during zero-shot evaluations. For open-domain QA, we don’t use any task specific instruction.
Task:Prompt used by “CoCoA”
[…]
Table 8: The prompt used by “CoCoA”.
Task Task Instruction
[…]
Refer to the information from the above two sources, verify the accuracy of the facts and the consistency of the logic, and choose the best prediction.
[…]
Table 9: A list of prompts used by CoCoA-zero.
Appendix G Limitations
While CoCoA has demonstrated excellent performance and provided valuable insights into collaboration with parametric and retrieved knowledge, there are still some limitations:
  • The current design focuses on a specific agent collaboration pattern via long-chain training. Its applicability to broader or alternative multi-agent architectures remains to be examined.
  • Although the approach performs robustly under limited supervision, its scaling with respect to larger models and datasets has not been systematically explored.
  • Although the performance has been improved, the token consumption has increased, which has certain limitations in practical applications. How to accelerate reasoning is still a future research direction.